The perimeter of a rectangle is 62ft. Describe the possible lengths of a side if the area of the rectangle is not to exceed 184 square feet
Answer by Boreal(15235)

P=2L+2W
half P=L+W=31 feet
L=31-W
area is W(31-W)=184
So -W^2+31W-184=0
W^2-31W+184=0
(W-23)(W-8)=0
If we make the longer side the length, L,
Length=23 ft, width=8 ft will give an area of 184 ft^2
Greater than or equal to 23 feet for length, and therefore <8 feet for width will give smaller values right up to 30 and 1
Length >=23 feet, 23+x by 8-x, where x is [0,8)
Width <= 8 feet